June wants to have $2,000,000 for her retirement in 30 years. How much should she save annually if she thinks she can earn 10% on her investments?
Answer #1

June save money at the end of every year

So,future value of annuity =P((1+r)^n -1)/r

$2,000,000=P((1.10)^30 -1)/0.10

P=2,000,000×0.10 /((1.10)^30 -1)

P= $12,158.79

Where P i's annual savings

r =rate of interest

n =number of years